# A Survey of 3D Reconstruction with Event Cameras: From Event-based Geometry to Neural 3D Rendering

### v1

@@ 4.4 Methods with 3D Gaussian Splatting
End-to-end learning: Some methods focus on directly learning the final outputs from end-to-end learning structure. In 2024, Wang et al. [93] proposed EvGGS. The model connects depth estimation, intensity reconstruction, and 3D Gaussian parameter regression in a collaborative learning framework. The joint training improves reconstruction accuracy and rendering efficiency. In the same year, Wu et al. [97] proposed Ev-GS. The method combines computational neuromorphic imaging and 3D Gaussian Splatting. It models logarithmic brightness changes and achieves fast convergence with purely event-based supervision.
5 Multimodal Methods with Event Cameras
Figure 10: Overview of multimodal methods with traditional outputs. (a) Schematic of multi-view event cameras working with structured light projector from Xiao et al. [99] (b) Schematic of a stereo setup of a D-RGB camera and a monocular event camera from Zuo et al. [121]
@@ 5 Multimodal Methods with Event Cameras
Multimodal 3D reconstruction refers to methods that combine event camera data with other sensing modalities, such as RGB images, depth sensors, or structured light systems. These methods aim to enhance reconstruction performance by leveraging the complementary strengths of each modality: the high temporal resolution and low latency of event cameras, and the rich spatial or appearance information provided by other sensors. Compared to stereo and monocular setups, multimodal systems are more flexible in input configurations. They often achieve higher reconstruction accuracy and robustness in low-light, high-speed, or motion-blurred environments.
@@
Early research in this direction typically combines event cameras with active sensing devices like structured light projectors [49], enabling high-speed depth recovery through time-coded patterns. Many recent works fuse asynchronous events with frame-based RGB or D-RGB data [121] to construct coloured 3D models, or integrate them into advanced neural rendering frameworks such as Neural Radiance Fields and 3D Gaussian Splatting. These approaches enable photorealistic, temporally-aware 3D reconstruction under challenging real-world conditions.
@@
Based on the output representation and modelling strategy, multimodal methods can be broadly classified into three categories: (5.1) methods with traditional 3D outputs such as point clouds, (5.2) methods with Neural Radiance Fields, and (5.3) methods with 3D Gaussian Splatting.
5.1 Multimodal Methods with Traditional Outputs
@@
Combine with Structured light: Structured light is an active 3D sensing technique that projects coded patterns onto a surface and reconstructs depth via triangulation. When integrated with event cameras, Structured light systems can achieve robust, high-speed depth sensing under challenging conditions. Recent works have introduced several types of encoding and fusion strategies to use both spatial and temporal features of event streams. Table 7 and Figure 10 provide an overview of these methods. As most of them rely on self-collected datasets, datasets are not included in Table 7. Based on the type of complementary modality integrated with event data, we categorise these methods into the following two types:
@@
Leroux et al. [49] proposed one of the earliest event-based structured light systems by projecting frequency-tagged light patterns onto the scene. Each spatial region is illuminated with a unique modulation frequency, and the event camera decodes depth by associating detected frequencies with pixel locations. Huang et al. [36] combined structured light projection with digital image correlation (DIC) [105, 104] for high-speed scanning. Xiao et al. [99] employed alternating binary speckle patterns and DIC-based stereo matching for fast and accurate reconstruction. Fu et al. [20] introduced spatio-temporal coding (STC) with an enhanced matching scheme for improved stereo robustness. Li et al. [50] proposed eFPSL, using time-frequency analysis to extract high-SNR fringe maps from events and an event-count-based shadow mask to reduce errors.
@@
Combine with D-RGB: Another method fuses event data with depth-RGB (D-RGB) sensors for improved scene understanding. Zuo et al. [121] proposed DEVO, combining time surface maps from events and depth supervision from a calibrated sensor. Their system performs semi-dense 3D-2D edge alignment to estimate poses and incrementally build point clouds under fast motion and poor lighting.
5.2 Multimodal Methods with Neural Radiance Fields
@@
Multimodal Event-based NeRF methods combine event streams with RGB images to achieve high-quality coloured 3D reconstruction, leveraging the temporal precision of events and the rich appearance information of RGB frames to enhance structural accuracy and visual fidelity under motion blur and low-light conditions. Table 8 and Figure 11 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
Static scene reconstruction: Some methods focus on static scene reconstruction by integrating event supervision with blurry RGB images [69, 8, 70, 89], or by designing physically inspired mechanisms [71, 14] to improve NeRF training stability and output sharpness. In 2023, Qi et al. [69] proposed E²NeRF, which introduces a blur rendering loss and an event rendering loss to improve NeRF training under motion blur and low-light settings. In 2024, Cannici et al. [8] proposed Ev-DeblurNeRF, which uses the Event Double Integral (EDI) [64] and a learnable event camera response function (eCRF) to reconstruct sharp NeRFs. It performs well under severe motion blur. Later that year, Qi et al. [70] proposed E3NeRF, which combines blurry images and events with spatial-temporal attention. It introduces an event-enhanced rendering loss to guide learning and improves robustness in non-uniform motion and low-light scenes. In the same year, Tang et al. [89] proposed LSE-NeRF, which models sensor response differences with per-time embeddings and event-based reflectance mapping to recover high-quality NeRFs without strict alignment. Around the same time, Qi et al. [71] proposed EBAD-NeRF, which introduces event-driven bundle adjustment. It jointly optimises NeRF and camera poses using an intensity-change event loss and a photometric blur loss. Also in 2024, Chen et al. [14] proposed Event-ID, which is the first event-driven framework for intrinsic decomposition. It combines an event-based reflectance model and a multi-view strategy to recover geometry, materials, and lighting under extreme conditions.
@@
5.3 Multimodal Methods with 3D Gaussian Splatting
@@
Multimodal Event-based Gaussian Splatting methods achieve high-quality 3D reconstruction by combining event streams with RGB images. These multimodal approaches leverage the high temporal resolution of events and the rich appearance information of RGB frames to improve geometric accuracy and colour fidelity. Table 9 and Figure 12 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
Static scene reconstruction: Some methods focus on static scene reconstruction by using event streams to compensate for motion blur and improve 3D reconstruction from blurry images. In 2024, Yu et al. [108] proposed EvaGaussians, which introduces learnable camera pose offsets and jointly optimise blurry image trajectories and 3D Gaussians. It improves reconstruction accuracy under severe motion blur conditions. Later that year, Weng et al. [96] proposed EaDeblur-GS, incorporating an Adaptive Deviation Estimator (ADE) [13, 48] network and blur-specific losses. It enables real-time and high-fidelity 3D reconstruction from extremely blurred inputs. In the same year, Deguchi et al. [15] proposed E2GS, which combines Event-based Double Integral (EDI) [65] and event rendering loss to enhance blurry image recovery and achieve 3D Gaussian reconstruction.
@@
Dynamic scene reconstruction: Some methods achieve dynamic 3D reconstruction either by modelling continuous time supervision from event streams [101, 51] or by introducing non-rigid deformation fields to explicitly represent scene dynamics [103]. In 2024, Xiong et al. [101] proposed Event3DGS, which introduces sparsity-aware sampling and a progressive training strategy. It reconstructs geometrically consistent and efficient 3D structures from event streams under high-speed egomotion. In the same year, Liao et al. [51] proposed EF-3DGS, which is the first event-aided 3DGS framework supporting free-trajectory video reconstruction. It enhances reconstruction accuracy and robustness in dynamic scenes by combining a Linear Event Generation Model (LEGM) [22, 25, 113] with a contrast maximisation-based image sharpening strategy [23, 31, 67, 88]. Also in 2024, Xu et al. [103] proposed Event-Boosted Deformable 3D Gaussians, which is the first deformable 3DGS framework incorporating event cameras. It applies a GS-threshold joint modelling strategy and a dynamic-static decomposition mechanism to achieve high-quality and efficient dynamic reconstruction.
6 Datasets
Table 10: Publicly available datasets for 3D reconstruction tasks with event cameras.
Abbr.: E = event stream, RGB = image frames, Li = LiDAR, IMU = inertial unit, Vcn = Vicon.
@@ 6 Datasets
SynthEVox3D[12] ICVR (2023) Synthetic Blender Simulator Voxel Grid 32 GB
@@
EventNeRF[43] CVPR (2023) Synthetic Mono E + RGB Refs RGB Frame 18 GB
DAVIS240C[60] IJRR (2017) Hybrid Mono 240$\times$180 E, IMU Pose 10 GB
Although numerous event-based vision datasets have been introduced, only a limited subset is suitable for 3D reconstruction, and many of them are not specially designed for 3D reconstruction. Many recent works rely on private datasets (e.g., [96]) or synthesise events from RGB videos using simulators without releasing the resulting data (e.g., [108], [103], [87]). As a result, only a small number of publicly available datasets provide the dense depth, stereo disparity, or sub-millimetre trajectories required for rigorous evaluation of reconstruction performance. Table 10 summarises the main publicly available datasets, and Table 11 provides example visualisation of sensors, RGB frames, event frames, and labels from the representative datasets. These datasets naturally fall into four categories based on their data acquisition setting, type of geometric supervision, and target reconstruction tasks: (1) outdoor navigation datasets with LiDAR or visual-inertial ground truth, (2) indoor object and human reconstruction datasets with motion capture or laser scans, (3) synthetic datasets with dense annotations, and (4) photometric benchmarks with accurate camera poses but without metric depth.
@@
Table 11: Representative Datasets used for event-based 3D Tasks.
Dataset Sensors Frames Events Labels DSEC  [27][28] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] MVSEC  [120] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] TUM-VIE  [42] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] IMU Data ViViD++  [47] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EV-IMO2  [6] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SynthEVox3D  [12] Synthetic Dataset [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EventScape  [26] Synthetic Dataset [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SEVD  [1] Synthetic Dataset [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
7 Research Gaps & Future Direction

### v2

@@ IV-D Methods with 3D Gaussian Splatting
End-to-end learning: Some methods focus on directly learning the final outputs from end-to-end learning structure. In 2024, Wang et al. [131] proposed EvGGS that connects depth estimation, intensity reconstruction, and 3D Gaussian parameter regression in a collaborative learning framework. The joint training improves reconstruction accuracy and rendering efficiency. In the same year, Wu et al. [133] proposed Ev-GS that combines neuromorphic imaging with 3DGS, modelling logarithmic brightness changes and enabling fast convergence using only event-based supervision.
V Multimodal Methods with Event Cameras
@@
Multimodal 3D reconstruction refers to methods that combine event camera data with other sensing modalities [139, 140, 141]. These methods aim to enhance reconstruction performance by leveraging the complementary strengths of each modality: the high temporal resolution and low latency of event cameras, and the rich spatial or appearance information provided by other sensors. Compared to stereo and monocular setups, multimodal systems are more flexible in input configurations. They often achieve higher reconstruction accuracy and robustness in low-light, high-speed, or motion-blurred environments.
@@
Early research in this direction typically combines event cameras with active sensing devices, such as structured light projectors [34], which enable high-speed depth recovery through time-coded patterns, and RGB-depth (RGB-D) cameras [140], which provide readily available depth data for event fusion. Many recent works fuse asynchronous events with frame-based RGB to construct coloured 3D models, or integrate them into advanced neural rendering frameworks such as Neural Radiance Fields and 3D Gaussian Splatting [139, 142]. These approaches enable photorealistic, temporally-aware 3D reconstruction under challenging real-world conditions.
@@
Based on the output representation and modelling strategy, multimodal methods can be broadly classified into three categories: (V-A) methods with traditional 3D outputs such as point clouds, (V-B) methods with Neural Radiance Fields, and (V-C) methods with 3D Gaussian Splatting.
V-A Multimodal Methods with Traditional Outputs
@@
Multimodal methods with traditional outputs recover geometric structures by combining event data with complementary sensors such as structured light or RGB-D cameras. These systems fuse the temporal precision of events with the spatial density of external inputs, enabling robust reconstruction under challenging conditions. Table VII and Figure 13 provide an overview of these methods. As most of them rely on self-collected datasets, datasets are not included in Table VII. Based on the type of complementary modality integrated with event data, we categorise these methods into the following two types:
Combine with Structured light: Structured light is an active 3D sensing technique that projects coded patterns onto a surface and reconstructs depth via triangulation [143]. When integrated with event cameras, structured light systems can achieve robust, high-speed depth sensing under challenging conditions. Recent works have introduced several types of encoding and fusion strategies to use both spatial and temporal features of event streams.
@@
In 2015, Matsuda et al. [141] proposed MC3D, one of the earliest event-based structured light systems. The method correlates laser scan timing with event timestamps, achieving high-precision depth reconstruction with per-pixel single-shot efficiency. In 2018, Leroux et al. [34] proposed event-based structured light systems by projecting frequency-tagged light patterns onto the scene. Each spatial region is illuminated with a unique modulation frequency, and the event camera decodes depth by associating detected frequencies with pixel locations. In 2021, Huang et al. [144] combined structured light projection with digital image correlation (DIC) [145, 146] for high-speed scanning. Also in 2021, Muglikar et al. [147] proposed ESL, which estimates depth by maximising spatio-temporal consistency between a laser projector and an event camera. By processing events in local space-time regions, their method improves robustness to noise. ESL accurately estimates depth in challenging scenes but is not real-time due to high computational cost. In 2023, Xiao et al. [148] employed alternating binary speckle patterns and DIC-based stereo matching for fast and accurate reconstruction. Also in 2023, Fu et al. [149] introduced spatio-temporal coding (STC) with an enhanced matching scheme for improved stereo robustness. In 2024, Li et al. [150] proposed eFPSL, using time-frequency analysis to extract high-SNR fringe maps from events and an event-count-based shadow mask to reduce errors.
Figure 13: Overview of multimodal methods with traditional outputs. (a) Schematic of event cameras working with structured light projector from Xiao et al. [148]. (b) Schematic of a stereo setup of an RGB-D camera and a monocular event camera from Zuo et al. [151].
TABLE VII: Event cameras in multimodal system: Methods with traditional outputs
Author Year Inputs Priors Event Rep. Real-time Output (Dense?)
Weikersdorfer et al. [140] 2014 RGB-D, Event stream Trajectory Event-by-event Point cloud (✗)
Matsuda et al. [141] 2015 Structured light, Event stream - Event-by-event 3D depth map (✓)
Leroux et al. [34] 2018 Structured light, Event stream Pose Time surface Point cloud (✓)
Huang et al. [144] 2021 Structured light, Event stream Pose Event-by-event Point cloud (✓)
Muglikar et al. [147] 2021 Structured light, Event stream - Event-by-event 3D depth map (✓)
Zuo et al. [151] 2022 RGB-D, Event stream Trajectory Time surface Point cloud (✗)
Xiao et al. [148] 2023 Structured light, Event stream Pose Event frame Point cloud (✓)
Fu et al. [149] 2023 Structured light, Event stream Pose Time surface 3D depth map (✗)
Li et al. [150] 2024 Structured light, Event stream Pose Event-by-event 3D depth map (✗)
@@
Combine with RGB-D: Some methods fuse event data with RGB-D sensors for improved scene understanding from depth and RGB information. In 2014, Weikersdorfer et al. [140] proposed EB-SLAM-3D, a novel event-based 3D SLAM algorithm using a D-eDVS, enabling low-power, low-latency mapping with a sparse voxel grid at 20× real-time speed. In 2022, Zuo et al. [151] proposed DEVO, combining time surface maps from events and depth supervision from a calibrated sensor. Their system performs semi-dense 3D-2D edge alignment to estimate poses and incrementally build point clouds under fast motion and poor lighting.
V-B Multimodal Methods with Neural Radiance Fields
@@
Multimodal Event-based NeRF methods combine event streams with RGB images to achieve high-quality coloured 3D reconstruction, leveraging the temporal precision of events and the rich appearance information of RGB frames to enhance structural accuracy and visual fidelity under motion blur and low-light conditions. Table VIII and Figure 14 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
Static scene reconstruction: Some methods focus on static scene reconstruction by integrating event supervision with blurry RGB images [152, 154, 155, 156], or by designing physically inspired mechanisms [157, 158] to improve NeRF training stability and output sharpness. In 2023, Qi et al. [152] proposed E²NeRF, which introduces a blur rendering loss and an event rendering loss to improve NeRF training under motion blur and low-light settings. In 2024, Cannici et al. [154] proposed Ev-DeblurNeRF, which uses the Event Double Integral (EDI) [159] and a learnable event camera response function (eCRF) to reconstruct sharp NeRFs. It performs well under severe motion blur. Later that year, Qi et al. [155] proposed E3NeRF, which combines blurry images and events with spatial-temporal attention [119]. It introduces an event-enhanced rendering loss to guide learning and improves robustness in non-uniform motion and low-light scenes. In the same year, Tang et al. [156] proposed LSE-NeRF, which models sensor response differences with per-time embeddings and event-based reflectance mapping to recover high-quality NeRFs without strict alignment. Later, Li et al. [160] proposed BeNeRF, which jointly reconstructs photorealistic 3D scenes and estimates camera motion from a single blurry RGB image and event stream, using an event rendering loss and B-spline trajectory representation for end-to-end optimisation. Around the same time, Qi et al. [157] proposed EBAD-NeRF, which introduces event-driven bundle adjustment. It jointly optimises NeRF and camera poses using an intensity-change event loss and a photometric blur loss. Also in 2024, Chen et al. [158] proposed Event-ID, which is the first event-driven framework for intrinsic decomposition. It combines an event-based reflectance model and a multi-view strategy to recover geometry, materials, and lighting under extreme conditions. In 2025, Chen et al. [161] proposed EvHDR-NeRF, which models a radiance-based relationship that accounts for exposure time and the camera response function (CRF), enabling HDR 3D reconstruction from single-exposure LDR images and event streams.
TABLE VIII: Event cameras in multimodal system: NeRF-based methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Qi et al. [152] E2NeRF 2023-10 Blurry RGB, Event stream EAF Self-collected [152]
Ma et al. [139] Deformable Event-NeRF 2023-10 Blurry RGB, Event Stream Event-by-event HS-ERGB [162]
Cannici et al. [154] Ev-DeblurNeRF 2024-06 Blurry RGB, Event stream EAF Ev-DeblurBlender [163]
Qi et al. [155] E3NeRF 2024-08 Blurry RGB, Event stream EGA Self-collected [155]
Tang et al. [156] LSE-NeRF 2024-09 Blurry RGB, Event stream EAF Self-collected [156], EVIMOv2 [164]
Li et al. [160] BeNeRF 2024-10 Blurry RGB, Event stream EAF Qi et al. [152]
Qi et al. [157] EBAD-NeRF 2024-10 Blurry RGB, Event stream EAF Self-collected [157]
Chen et al. [158] Event-ID 2024-10 Blurry RGB, Event stream EAF Self-collected [158]
Rudnev et al. [153] Dynamic EventNeRF 2024-12 Blurry RGB, Event stream EAF Self-collected [153]
Chen et al. [161] EvHDR-NeRF 2025-04 LDR RGB, Event stream EAF HDR-NeRF dataset [165]
@@
V-C Multimodal Methods with 3D Gaussian Splatting
@@
TABLE IX: Event cameras in multimodal system: 3DGS-based methods
Author Model Yr-Mo Inputs Event Rep. Colourful Dataset
Yu et al. [142] EvaGaussians 2024-05 Blurry RGB, Event stream EGA Self-collected [142]
Xiong et al. [167] Event3DGS 2024-06 Blurry RGB, Event stream EGA Self-collected [167]
Weng et al. [55] EaDeblur-GS 2024-07 Blurry RGB, Event stream EGA Qi et al. [152]
Liao et al. [56] EF-3DGS 2024-10 Blurry RGB, Event stream EAF Tanks and Temples [168]
Deguchi et al. [29] E2GS 2024-10 Blurry RGB, Event stream EAF Self-collected [29]
Xu et al. [166] EventBoosted-3DGS 2024-11 Blurry RGB, Event stream EGA Self-collected [166]
Huang et al. [169] Ev3DGS 2024-12 Blurry RGB, Event stream EAF Qi et al. [152]
Wu et al. [170] SweepEvGS 2024-12 Static RGB, Event stream EGA Self-collected [170]
Lee et al. [171] Sensor Fusion Splatting 2025-02 RGB, Event stream, Depths EGA Self-collected [171]
Matta et al. [172] BeSplat 2025-03 Blurry RGB, Event stream EAF BeNeRF dataset [173]
Deng et al. [174] EBAD-Gaussian 2025-04 Blurry RGB, Event stream EAF Qi et al. [157]
Figure 16: Multimodal 3D Gaussian reconstruction based on blurry RGB and event data originates from the E2GS method [29] under CC BY 4.0. @IEEE 2024. The event modality assists in deblurring scene rendering.
Multimodal Event-based Gaussian Splatting methods achieve high-quality 3D reconstruction by combining event streams with RGB images. These multimodal approaches leverage the high temporal resolution of events and the rich appearance information of RGB frames to improve geometric accuracy and colour fidelity. Table IX and Figure 15 provide an overview of these methods. Based on the type of scene being reconstructed, we categorise these methods into the following two types:
Static scene reconstruction: Some methods focus on static scene reconstruction by using event streams to compensate for motion blur and improve 3D reconstruction from blurry images. In 2024, Yu et al. [142] proposed EvaGaussians, which introduces learnable camera pose offsets and jointly optimises blurry image trajectories and 3D Gaussians. It improves reconstruction accuracy under severe motion blur conditions. Later that year, Weng et al. [55] proposed EaDeblur-GS, incorporating an Adaptive Deviation Estimator (ADE) [175, 176] network and blur-specific losses. It enables real-time and high-fidelity 3D reconstruction from extremely blurred inputs. In the same year, Deguchi et al. [29] proposed E2GS, which combines Event-based Double Integral (EDI) [177] and event rendering loss to enhance blurry image recovery and achieve 3D Gaussian reconstruction. Also in 2024, Huang et al. [169] proposed Ev3DGS, which introduces blur and event rendering losses to guide 3D Gaussian Splatting, enabling fast and accurate colour reconstruction from blurry RGB images and event streams. In 2025, Matta et al. [172] proposed BeSplat, which reconstructs sharp and colour-consistent 3D scenes from a single blurry RGB image and event stream by jointly optimising scene representation and camera motion with event-guided spatio-temporal lifting. Later, Deng et al. [174] proposed EBAD-Gaussian, which leverages complementary image and event modalities and introduces event-driven bundle adjustment with motion blur modelling to jointly optimise Gaussians and camera poses, enabling sharp and physically consistent 3D reconstruction under severe motion blur. Figure 16 provides a visualisation of its result compared to RGB-based 3D Gaussian Splatting [69].
@@
Dynamic scene reconstruction: Some methods achieve dynamic 3D reconstruction either by modelling continuous time supervision from event streams [167, 56] or by introducing non-rigid deformation fields to explicitly represent scene dynamics [166, 171]. In 2024, Xiong et al. [167] proposed Event3DGS, which introduces sparsity-aware sampling and a progressive training strategy. It reconstructs geometrically consistent and efficient 3D structures from event streams under high-speed egomotion. In the same year, Liao et al. [56] proposed EF-3DGS, which is the first event-aided 3DGS framework that supports free-trajectory rendering in dynamic scenes. It enhances reconstruction accuracy and robustness in dynamic scenes by combining a Linear Event Generation Model (LEGM) [178, 179, 180] with a contrast maximisation-based image sharpening strategy [181, 182, 183, 184]. Also in 2024, Xu et al. [166] proposed Event-Boosted Deformable 3D Gaussians, which is the first deformable 3DGS framework incorporating event cameras. It applies a GS-threshold joint modelling strategy and a dynamic-static decomposition mechanism to achieve high-quality and efficient dynamic reconstruction. In 2025, Lee et al. [171] proposed Sensor Fusion Splatting, which fuses RGB images, event streams, and depth maps from an RGB-D camera with deformable Gaussians and modality-specific losses, enabling high-quality 3D reconstruction in dynamic scenes.
VI Datasets
Although numerous event-based vision datasets have been introduced, only a limited subset is suitable for 3D reconstruction, and many of them are not specially designed for 3D reconstruction. Many recent works rely on private datasets (e.g., [55]) or synthesise events from RGB videos using simulators without releasing the resulting data (e.g., [142], [166], [72]). As a result, only a small number of publicly available datasets provide the dense depth, stereo disparity, or sub-millimetre trajectories required for rigorous evaluation of reconstruction performance. Table X summarises the main publicly available datasets, and Table XI provides example visualisation of sensors, RGB frames, event frames, and labels from the representative datasets. These datasets naturally fall into four categories based on their data acquisition setting, type of geometric supervision, and target reconstruction tasks: (1) outdoor navigation datasets with LiDAR or visual-inertial ground truth, (2) indoor object and human reconstruction datasets with motion capture or laser scans, (3) synthetic datasets with dense annotations, and (4) photometric benchmarks with accurate camera poses but without metric depth.
TABLE X: Publicly available datasets for 3D reconstruction tasks with event cameras. (Abbr.: E = event stream, RGB = image frames, Li = LiDAR, IMU = inertial unit, Vcn = Vicon.)
@@ VI Datasets
EventNeRF [35] CVPR (2023) Synthetic Mono E + RGB Refs RGB Frame 18 GB
SynthEVox3D [95] ICVR (2023) Synthetic Mono E (E2V) from Blender Simulator Voxel Grid 32 GB
@@
DAVIS240C [101] IJRR (2017) Real + Synthetic Mono 240$\times$180 E, IMU Pose 10 GB
@@
TABLE XI: Representative datasets used for event-based 3D Tasks. All the synthetic datasets are created by Blender [189]. Parts of the figures in the table are referenced from Ghosh et al. with permission [190].
Dataset Sensors Frames Events Labels DSEC  [93][83] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] MVSEC  [80] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] TUM-VIE  [124] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] ViViD++  [186] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EV-IMO2  [164] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SynthEVox3D  [95] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EventScape  [187] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SEVD  [188] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
VII Research Gaps & Future Direction